# Moments and the Basic Equilibrium Equations

### v3

@@ Abstract.
We develop the classical theory of moments of configurations of weighted points with a focus on systems with an identically vanishing first moment. The latter condition produces equations for equilibrium configurations of systems of interacting particles under the sole condition that interactions are pairwise and along the line determined by each pair of particles. Complying external forces are admitted, so the description of some dynamical equilibrium configurations, such as relative equilibria in Celestial Mechanics, is included in our approach. Moments provide a unified framework for equilibrium problems in arbitrary dimensions. The equilibrium equations are homogeneous, of relatively low degree, invariant by oriented isometries (for interactions depending only on mutual distances), and are obtained through simple algebraic procedures requiring neither reduction by isometries nor a variational principle for their determination. Our equations include the renowned sets of $n$-body central configuration equations by O. Dziobek, and by A. Albouy and A. Chenciner. These equations are extended to a rather broad class of equilibrium problems, and several new sets of equilibrium equations are introduced. We also apply moments to establish a theory of constraints for mutual distances of configurations of fixed dimension and of co-spherical configurations, thus re-obtaining and adding to classical results by A. Cayley, among others.
@@
Affine Geometry, Barycentric Coordinates, Equilibria, $N$-Body Problem, Central Configurations.
@@ 1. Introduction
Equilibrium problems have contributed to the development of Mathematics and Physics since ancient Greece. The work of Archimedes on the principle of the lever is considered the beginning of the mathematical treatment of Statics, but even before him, there had appeared books attributed to Aristotle and to Euclid on the dynamics and balance of the lever, respectively [15]. The concept of barycenter, or center of gravity, of a system of point masses, and its application to studying equilibrium configurations go back to Archimedes. As we shall review in section 2, the barycenter is defined for sets of weighted points whose total weight, $\mu_{0}$, is nonzero as the unique point where the first moment, $\mu_{1}$, vanishes. In the 17th century, C. Huygens used $\mu_{1}$ to introduce the notion of expected value in probability, and realized the importance of the second moment, $\mu_{2}$, in the form of the moment of inertia (Euler’s terminology), for the analysis of pendulum motion. Later on G. W. Leibniz and J. S. König applied $\mu_{2}$ in Geometry and Dynamics [19]. A consequential event was the invention by A. F. Möbius of homogeneous, barycentric coordinates in the early 19th century [13]. In contemporary Mathematical Statistics and Analysis, among other areas, one finds the classical (Stieltjes, Hausdorff, Hamburger) moment problem: given a sequence of moments of a real random variable, determine the (unique) probability distribution (measure) that produced the sequence [27, 30].
@@
In the physical sciences, it has been a long tradition to represent configurations of point masses, electric charges, point vortices, etc, by a finite set of points $X$ in the Euclidean affine space $\mathbb{A}=\mathbb{R}^{N}$ over the field of scalars $K=\mathbb{R}$, and a function $w$ from $X$ to $K$ denoting the mass, charge, vorticity, etc, of each point of $X$. In an abstract sense, we refer to $X$ and $w$ as configuration and weight function, respectively, and call the pair $(X,w)$ a weighted system. Three elementary functions are associated with every weighted system: the total weight $\mu_{0}$ (“zeroth moment”), the first moment $\mu_{1}$ and the second moment $\mu_{2}$. In the case of a configuration of point masses, $\mu_{0}$ is the total mass and, besides the previously mentioned direct relation between the barycenter and $\mu_{1}$, and between the moment of inertia and $\mu_{2}$, if one allows the configuration to vary differentiably with time, then $\mu_{1}$ can be used to define the total linear momentum, while $\mu_{2}$ becomes (twice) the kinetic energy. All these concepts naturally extend to discrete and continuous settings.
@@
Drawing apart from tradition, in this paper we view the interactions of particles as defining the weight functions themselves. More precisely, we introduce a family of weighted systems $(X,w_{x})$, $x\in X$, associated with a configuration $X$ of interacting particles, so that $w_{x}(y)$ is the force (this familiar term will be often used in place of interaction) exerted by the particle $y\in X$ on the particle $x$, while $w_{x}(x)$ is the negative of the total force acting on $x$. This basic remark, and the key realization that the equilibrium configurations correspond to fulfilling the condition $\mu_{1}=\overrightarrow{O}$ (this notation means that the first moment is the null function) for each system $(X,w_{x})$, motivates a deeper investigation of the theory of moments of weighted systems in Euclidean affine spaces. At the same time, we have a bridge to applications of moments to problems of equilibrium configurations of interacting particles possibly subject to certain types of external forces. To be more explicit, all our equilibrium equations are valid for systems of interacting weighted points satisfying the single fundamental hypothesis:
(FH): The interaction between the particle at $x$ and the particle at $y$, with $x,y\in X$, is directed along the line determined by $x$ and $y$.
As a matter of fact, our formalism holds for systems satisfying the weaker hypothesis:
(WFH): The total interaction (e.g., internal plus external forces) at each point $x$ of the configuration $X$ is a linear combination of the relative position vectors from $x$ to the remaining points of $X$.
Hypothesis (WFH) extends the scope of applications of this work to systems affected by some inertial forces, for instance, centrifugal forces which give rise to relative equilibrium configurations.
We shall see that the theory of moments of weighted systems provides both a unified framework for the construction of families of equilibrium equations with distinctive properties, as well as a tool to study geometric aspects of the configurations themselves. Two important points of departure for our work were the classical identities deduced by the pioneers of moment theory, and the investigations on central configurations of the $n$-body problem.
@@
Central configurations is a research topic with a long history and recognized relevance to Celestial Mechanics [4, 3, 28, 31, 34]. One of the reasons is that the only known explicit solutions of the Newtonian $n$-body problem, namely the homographic motions, start at such configurations (cf. Wintner [34], §375). The possible homographic motions include the relative equilibria, which are planar periodic solutions, and, in recent times, have been shown to occur in any Euclidean space of even dimension [3, 28]. In the context of the $n$-body problem, it is common to use the terminologies planar (or two-dimensional) central configurations and relative equilibria interchangeably. Notwithstanding the many particular questions that have been answered, so far a classification of central configurations is known only for three arbitrary masses, thanks to Euler [17] and Lagrange [26], and for four equal masses, thanks to Albouy [2], who based his work on a previously obtained symmetry result by himself [1]. General, basic problems on central configurations remain unsolved. An outstanding example is the celebrated Smale’s sixth problem for the 21st century [32, 33], which asks whether the number of similarity classes of relative equilibria of $n$ positive point masses is finite for all $n$. In the broader sense of central configurations of arbitrary dimension, Smale’s problem is part of the Chazy-Wintner-Smale finiteness conjecture [4]. The finiteness problem has been fully solved only for $n=4$, by Hampton and Moeckel [22]. Considering just generic finiteness with respect to the masses, finiteness has been proved for $n=5$ in dimension two by Albouy and Kaloshin [5], and for the central configurations of $n$ bodies in dimension n2n-2, the so-called Dziobek configurations, by Moeckel [29] (see also Dias [14]). Below we consider the most important sets of central configuration equations used in the classification and finiteness problems, interspersed with summaries of our present contributions.
@@
In the seminal paper [16], O. Dziobek used mutual distances as coordinates to describe central configurations as critical points of a modified potential function subject to constraints that guaranteed the configuration was planar. The constraints used by Dziobek were expressed in terms of the determinants introduced by A. Cayley [10], which later became known as Cayley-Menger determinants [8, 9]. For the planar four-body problem, Dziobek produced relatively simple algebraic equations expressing the reciprocal cubed mutual distances $r_{ij}^{-3}$ as an affine function of the product of suitably defined (volume) barycentric coordinates, $\Delta_{i}\Delta_{j}$, divided by the product of the masses, $m_{i}m_{j}$. His approach was adopted and developed by Albouy in the previously mentioned classification of four-equal-mass central configurations [1, 2]. Later on, Moeckel [29] adapted Dziobek’s method to dimension n2n-2, and, using barycentric coordinates, Albouy [4] provided a simple algebraic path to the Moeckel-Dziobek equations which did not require the (typically large) Cayley-Menger determinant.
Notably the whole Dziobek formalism is contained in the theory of weighted systems $(X,w)$ with $\mu_{1}=\overrightarrow{O}$. The weight functions of such systems (with fixed $X$) form a vector space, $\mathbb{W}_{0}(X)$, and we describe a method for constructing bases for $\mathbb{W}_{0}(X)$ using volume barycentric coordinates. The case $\dim\mathbb{W}_{0}(X)=1$ of our theory, and the fundamental hypothesis (FC) of page 1, are implicit in Albouy’s algebraic approach: in essence, it is the ideas in reference [4] that we recast and generalize to $\dim\mathbb{W}_{0}(X)\geq 1$. Besides the intensive use of barycentric coordinates, this generalization requires some additional concepts, in particular that of codimension of a configuration, which plays a chief role throughout the remainder of this paper. A significant number of equilibrium (including central configuration) equations is introduced through the “codimension-exterior power formula”
(1.1) $\bigwedge^{c+1}\mathbb{W}_{0}(X)=\{O\},$
where c=nd1c=n-d-1 is the codimension of $X$, $n$ is the number of points in $X$, and $d$ is dimension of the affine closure of $X$, usually referred to as the dimension of $X$. It would not be an overstatement to say that (1.1) synthesizes the basic theory of systems with $\mu_{1}=\overrightarrow{O}$. Formula (1.1) is used in section 5 to provide a broad generalization of some determinantal identities obtained by Dziobek [16, 21] for central configurations with codimension one. These identities yield tailored equilibrium equations that exploit properties such as symmetry and dimension, as illustrated in example 5.3. In view of its versatility, we deem equation (1.1) one of the main contributions of our work. We also remark that the Dziobek-type homogeneous, quadratic central configuration equations of Dias [14] are here extended to arbitrary weighted systems with $\mu_{1}=\overrightarrow{O}$, and we show that they actually provide a characterization of such systems having any specific dimension. Finally, a somewhat surprising connection with Algebraic Geometry arises as we interpret barycentric coordinates as Plücker coordinates: Plücker relations [24, 25] for Grassmannians of $d$-dimensional subspaces of $\mathbb{R}^{n-1}$ emerge directly from the generalization of Dias’ formulæ.
Regarding applications, two remarks seem in order. Firstly, Plücker relations, the linear relations for volume barycentric coordinates which form our basis vectors of $\mathbb{W}_{0}(X)$, and the generalized Dziobek equations of section 6 may be used to formulate equilibrium problems entirely in terms of barycentric coordinates, via corollary 2.6. And secondly, all the weight functions defined by the models of interacting particles having $X$ as an equilibrium configuration coinhabit the space $\mathbb{W}_{0}(X)$. Thus they are a fortiori interrelated, and their relationships may be expressed through equations such as those produced by formula (1.1). In this context, our bases for $\mathbb{W}_{0}(X)$ supply a purely geometric (i.e. non-physical) common framework for representing the weight functions of each model individually.
Going back to the topic of central configurations, it is worth noting that Dziobek equations also found an application in the pioneering resolution of the finiteness problem for gravitational relative equilibrium configurations [22], as well as for stationary configurations of vortices [23]. However, the main formulation of these finiteness problems employed a different set of central configuration equations, namely the one introduced by A. Albouy and A. Chenciner in the celebrated article [3]. Briefly, the finiteness proofs in [22, 23] applied ideas from Algebraic Geometry to exclude the possibility of a continuum of solutions, taking advantage of the fact that the so-called Albouy-Chenciner equations are algebraic. A remarkable feature of the Albouy-Chenciner equations is that they depend only on the mutual distances of the configuration which, besides eliminating the rotational and translational symmetries, makes them at once valid for configurations of all dimensions.
Unlike the known deductions of the Albouy-Chenciner equations [3, 22], which use reduction by isometries, early on in the paper we obtain these equations for every system with $\mu_{1}=\overrightarrow{O}$ as a nearly straightforward consequence of the basic interplay of the moments $\mu_{0}$ and $\mu_{1}$ with the second moment $\mu_{2}$. Later on, we provide another set of algebraic equations exhibiting the properties of the Albouy-Chenciner equations mentioned in the previous paragraph. We call these new equations the extended Leibniz identities, as they are closely related to a well-known identity involving moments and mutual distances attributed to Leibniz. We stress that, just as for the generalized Dziobek approach, the applications of the generalized Albouy-Chenciner and the extended Leibniz identities include all equilibrium problems satisfying the weaker fundamental hypothesis (WFC) stated on page 1. Moreover, concerning Geometry, the interplay of the first three moments allows us to establish a theory of constraints for distances from the points of a configuration to an arbitrary point in the ambient space, and for mutual distances between the points of a configuration of prescribed dimension or lying on a sphere. In particular, Cayley-Menger determinantal conditions for the dimension are derived, and the number of independent Cayley-Menger constraints needed to ensure the configuration has a given dimension is expressed in terms of the configuration’s codimension $c$ simply as $c+1\choose 2$}. Our formula for the number of constraints generalizes the ones for dimensions 1, 2 and 3 in Wintner [34], $\S 357$. It underscores once again the importance of the codimension in the study of configurations of points in Euclidean affine spaces.
@@
Before moving on to a summary of specific contents, we make a brief note related to the moment problem mentioned in the first paragraph. We notice that the moment problem suggests an inverse of the equilibrium problem treated in the present work, viz., assuming the configuration $X$ is known, one would be asked to find the interactions $w_{x}$, $x\in X$, which make $X$ an equilibrium. For this problem, our equilibrium equations seem to offer relatively simple conditions to determine the $w_{x}$. In particular, the interactions can be determined using the linear systems provided by equations (2.2) or (4.3). If the interactions are symmetric, i.e., if Newton’s third law of motion (popularly “Action == (1)(-1) $\times$ Reaction”) holds, then, using the generalized Dziobek equations of section 6, the $w_{x}$ can be directly expressed in terms of barycentric coordinates (or mutual distances) and $c+1\choose 2$} parameters, where $c$ is the codimension of $X$.
@@
The fundamentals of the theory of moments of weighted systems are reviewed in section 2, culminating with the Huygens-Leibniz-König (HLK) theorem, which collects classical results relating $\mu_{0},\mu_{1}$ and $\mu_{2}$. We demonstrate three corollaries of the HLK theorem with conditions on the second moment $\mu_{2}$ of a weighted system with $\mu_{0}=0$ to have $\mu_{1}=\overrightarrow{O}$. Prototypes of the generalized Albouy-Chenciner equations and the extended Leibinz identities appear at this point. Together with barycentric coordinates, which are introduced in section 4, the HLK theorem and its corollaries are the main tools for our formulations of equilibrium problems.
@@
Equilibrium equations for systems of interacting particles are firstly discussed in section 3. According to our fundamental hypothesis (FC), the total force on each particle $x\in X$ is a linear combinations with coefficients $\varphi_{x,y}$ of the relative positions from $x$ to the remaining particles $y\in X\setminus\{x\}$. We apply the results of section 2 to obtain equilibrium equations that are either linear or bilinear on the $\varphi_{x,y}$, and linear on the squared mutual distances between points of $X$. These equations contain, in particular, the generalization of the Albouy-Chenciner equations and the extended Leibniz identities.
Section 4 is dedicated to the set of all weighted systems with a fixed configuration $X$ and $\mu_{1}=\overrightarrow{O}$. The corresponding weight functions form a vector space $\mathbb{W}_{0}(X)$ whose dimension is the codimension of $X$. Equation (1.1) immediately follows. In section 4, we view Dziobek configurations as simplexes $S$ with an additional point in the affine closure of $S$. This geometric interpretation allow us to introduce devices called Dziobek trees for $X$ in order to construct explicit bases for $\mathbb{W}_{0}(X)$. Section 4 closes with a set of homogeneous second degree equations that characterize the weight functions in $\mathbb{W}_{0}(X)$. Remark 4.10 discusses applications to Grassmannians.
Section 5 concerns some illustrative applications of the previous sections to central configurations. After providing a formal definition of central configurations followed by a short review of basic results, some classical formulæ  are reobtained, and new central configuration equations are discussed. We consider the not so explored problem of central configurations with zero total mass, for which some geometric properties and algebraic relations are directly determined. In addition, we examine the planar five-body central configurations having three collinear bodies, assuming the total mass is nonzero. Our results impose some significant restrictions on the possible central configurations which are consistent with previous studies of particular cases by other authors.
Section 6 focuses on generalizing the set of equations found by Dziobek [16] to describe central configurations as constrained critical points. Dziobek’s equations assume their simplest form when $c=1$, and, in section 6, using matrices, we extend Albouy’s arguments [4] to any symmetric family of $n$ weight vectors in $\mathbb{W}_{0}(X)$ on a common configuration $X$ of arbitrary dimension. For systems of interacting particles, symmetry means the interactions obey Newton’s third law of motion, which, in the notation of section 3, translates simply as $\varphi_{x,y}=\varphi_{y,x}$, for all $x,y\in X$.
The article concludes with two appendices. In appendix A we show how the HLK theorem of section 2 can be used to find constraints for mutual distances. We apply corollary 2.6 to obtain constraints corresponding to either fixing the dimension of the configuration $X$, which take the form of vanishing Cayley-Menger determinants of Dziobek subconfigurations of $X$, or imposing co-sphericity, which are expressed as the vanishing of the determinant of the so-called relative configuration matrix [3, 28] of $X$. Two theorems are proved in appendix A. In theorem A.2.1 we clarify the relationship between the space of weight functions with $\mu_{1}=\overrightarrow{O}$, $\mathbb{W}_{0}(X)$, and the kernels of the relative configuration matrix and the Cayley-Menger matrix of $X$. The dimension of the configuration is thus related to vanishing Cayley-Menger determinants of subconfigurations. Then, in theorem A.2.2, we obtain the expression $c+1\choose 2$} for the number of independent constraints that ensures a configuration has a fixed dimension. Finally, in appendix B, we determine the kernel of the relative configuration matrix viewed, as in reference [3], as a bilinear form. The extended Leibniz identities follow therefrom, and a final set of central configuration formulæ  is provided.

### v4

# Moments, Equilibrium Equations and Mutual Distances

@@ Abstract.
We review and develop the classical theory of moments of configurations of weighted points with a focus on systems with an identically vanishing first moment. The latter condition produces equations for equilibrium configurations of systems of interacting particles under the sole condition that interactions are between pairs of particles and along the lines connecting such pairs. Complying external forces are admitted, so the description of some dynamical equilibrium configurations, such as relative equilibria in Celestial Mechanics, is included in our approach. Moments provide a unified framework for equilibrium problems in arbitrary dimensions. The equilibrium equations are homogeneous and invariant by isometries (for interactions depending only on mutual distances), and are obtained through simple algebraic procedures requiring neither reduction by isometries nor a variational principle for their determination. Our equations include the renowned set of $n$-body central configuration equations by A. Albouy and A. Chenciner. These equations are extended to a rather broad class of equilibrium problems, and new equilibrium equations written in terms of mutual distances are introduced. We also apply moments to the theory of constraints for mutual distances of configurations of fixed dimension, and for co-spherical configurations, thus re-obtaining and adding to classical results by A. Cayley and successors. For the sake of concreteness, novel sets of central configurations equations are provided.
@@
Affine Geometry, Equilibria, Mutual Distances, $N$-Body Problems, Central Configurations.
@@ 1. Introduction
Equilibrium problems have contributed to the development of Mathematics and Physics since ancient Greece. The work of Archimedes on the principle of the lever is considered the beginning of the mathematical treatment of Statics, but even before him, there had appeared books attributed to Aristotle and to Euclid on the dynamics and balance of the lever, respectively [12]. The concept of barycenter, or center of gravity, of a system of point masses, and its application to studying equilibrium configurations go back to Archimedes. As we shall review in section 2, the barycenter is defined for sets of weighted points whose total weight, $\mu_{0}$, is nonzero, as the unique point where the first moment, $\mu_{1}$, vanishes. In the 17th century, C. Huygens used $\mu_{1}$ to introduce the notion of expected value in probability, and realized the importance of the second moment, $\mu_{2}$, in the form of the moment of inertia (Euler’s terminology), for the analysis of pendulum motion. Later on G. W. Leibniz and J. S. König applied $\mu_{2}$ in Geometry and Dynamics [15]. A consequential event was the invention by A. F. Möbius of homogeneous, barycentric coordinates in the early 19th century [10]. In contemporary Mathematical Statistics and Analysis, among other areas, one finds the celebrated (Stieltjes, Hausdorff, Hamburger) moment problem: given a sequence of moments of a real random variable, determine the (unique) probability distribution (measure) that produced the sequence [20, 23].
@@
In the physical sciences, it has been a long tradition to represent configurations of (interacting) point masses, electric charges, point vortices, etc, by a finite set $X$ formed by $n$ points in the Euclidean affine space $\mathbb{A}=\mathbb{R}^{N}$ over the field of scalars $K=\mathbb{R}$, and a function $w$ from $X$ to $K$ denoting the mass, charge, vorticity, etc, of each point of $X$. In an abstract sense, we refer to $X$ and $w$ as configuration and weight function, respectively, and call the pair $(X,w)$ a weighted system. Moving beyond this tradition, in this paper we view the interactions of particles as themselves defining the weight functions. More precisely, we introduce a family of weighted systems $(X,w_{x})$, $x\in X$, associated with a configuration $X$ of interacting particles, so that $w_{x}(y)$ is the force (this familiar term will be often used in place of interaction) exerted by the particle $y\in X$ on the particle $x$, while $w_{x}(x)$ is the negative of the total force acting on $x$.
@@
As already mentioned, three elementary functions are historically associated with every weighted system: the total weight $\mu_{0}$ (“zeroth moment”), the first moment $\mu_{1}$ and the second moment $\mu_{2}$. The key realization that an equilibrium configuration $X$ corresponds to the family of weighted systems $(X,w_{x})$, $x\in X$, fulfilling the condition $\mu_{1}=\overrightarrow{O}$ (this notation means that the first moment is the null function) for each system $(X,w_{x})$ motivates a deeper investigation of the theory of moments of weighted systems in affine spaces. Moments allow us to formulate problems of equilibrium configurations of systems of interacting particles (possibly subject to certain types of external forces) that satisfy the single fundamental hypothesis (FH): the interaction between the particle at $x$ and the particle at $y$, with $x,y\in X$, is directed along the line determined by $x$ and $y$. Actually it suffices that the interactions fulfill the weaker fundamental hypothesis (WFH): the total interaction (e.g., the sum of all internal and external forces) at each point $x$ of the configuration $X$ is a linear combination of the relative position vectors from $x$ to the remaining points of $X$. Hypothesis (WFH) extends the scope of applications of the moment apparatus to systems affected by some inertial forces, for instance, centrifugal forces giving rise to relative equilibrium configurations.
The main pillar for the present article are the classical identities deduced by the pioneers of moment theory and their use, explicitly or not, to study central configurations of the $n$-body problem.
@@
Central configurations is a research topic with a long history and recognized relevance to Celestial Mechanics [2, 3, 21, 24, 27]. One of the reasons is that the only known explicit solutions of the Newtonian $n$-body problem, namely the homographic motions, start at such configurations (cf. Wintner [27], §375). The possible homographic motions include the relative equilibria, which are periodic solutions that occur in any Euclidean space of even dimension [2, 21]. Notwithstanding the many particular questions that have been answered, so far a classification of central configurations is known only for three arbitrary masses, thanks to Euler [14] and Lagrange [19], and for four equal masses, thanks to Albouy [1]. General, basic problems on central configurations remain unsolved. An outstanding example is the celebrated Smale’s sixth problem for the 21st century [25, 26], which asks whether the number of similarity classes of relative equilibria of $n$ positive point masses is finite for all $n$. In the broader sense of central configurations of arbitrary dimension, Smale’s problem is part of the Chazy-Wintner-Smale finiteness conjecture [3].
The finiteness problem for central configurations has been solved only in the case $n=4$ by Hampton and Moeckel [17, 18], who mainly employed the set of polynomial equations introduced by A. Albouy and A. Chenciner in the celebrated article [2]. A remarkable feature of the Albouy-Chenciner equations is that they depend only on the mutual distances of the configuration. Thus, besides eliminating the rotational and translational symmetries, the Albouy-Chenciner equations are at once valid for configurations of all dimensions. Unlike the known deductions of the Albouy-Chenciner equations [2, 17], which use reduction by isometries, early on in this paper we obtain a prototype of these equations for every system with $\mu_{1}=\overrightarrow{O}$ as a nearly straightforward consequence of the basic interplay of the moments $\mu_{0}$, $\mu_{1}$ and $\mu_{2}$. Taking advantage of the broader perspective offered by the theory of moments, we also provide another set of algebraic equations exhibiting the properties of the Albouy-Chenciner equations mentioned in the previous paragraph. We call these new equations the extended Leibniz identities, as they are closely related to a classical identity involving moments and mutual distances attributed to Leibniz.
In addition to providing a unified framework for the construction of families of equilibrium equations, we shall see that the theory of moments of weighted systems is a key tool to study the geometry of configurations.
@@
In the previously mentioned classification of four-equal-mass central configurations, Albouy adopted and developed the results by O. Dziobek [13], who used mutual distances as coordinates in order to describe $n$-body central configurations as critical points of a modified potential function subject to constraints that guaranteed the configuration was planar. The constraints used by Dziobek were expressed in terms of the determinants introduced by A. Cayley [8], which later became known as Cayley-Menger determinants [6, 7]. Dziobek’s apparatus was adapted by Moeckel [22] to dimension n2n-2, for arbitrary $n\geq 4$. The critical point approach to central configurations motivated us to revisit the theory of constraints for mutual distances. The interplay of the first three moments allows us to establish, in a simple manner, the theory of constraints for mutual distances between the points of a configuration of prescribed dimension, or lying on a sphere.
@@
Before moving on to a summary of each section in the paper, we make a brief note related to the moment problem mentioned in the first paragraph. We notice that the moment problem suggests an inverse of the equilibrium problem treated in the present work, viz., assuming the configuration $X$ is known, one would be asked to find the interactions $w_{x}$, $x\in X$, which make $X$ an equilibrium. For this problem, our equilibrium equations offer low-degree homogeneous relations that can be used to determine the $w_{x}$.
@@
The fundamentals of the theory of moments of weighted systems are reviewed in section 2, culminating with the Huygens-Leibniz-König (HLK) theorem, which collects classical results relating $\mu_{0},\mu_{1}$ and $\mu_{2}$. We demonstrate three consequential corollaries of the HLK theorem with conditions on the second moment $\mu_{2}$ of a weighted system with $\mu_{0}=0$ to have $\mu_{1}=\overrightarrow{O}$.
@@
Equilibrium equations for systems of interacting particles are introduced in section 3. We apply the results of section 2 and section 4 to obtain equilibrium equations that are either linear or bilinear on the $\varphi_{x,y}$, and linear on the squared mutual distances between points of $X$. These equations are the generalization of the Albouy-Chenciner equations for central configurations as well as the set of equations we call the extended Leibniz identities.
In section 4 we use matrices to provide a more algebraic approach to the study of moments, which are viewed as linear maps on the space of weight functions. The kernels of these maps are related in a way precisely stated in theorem 4.1, which serves as basis for the study of mutual distance constraints carried out in section 5. The section closes with a proposition that justifies the extended Leibniz identities stated in section 3.
The study of weighted systems $(X,w)$ with $\mu_{1}=\overrightarrow{O}$ calls for some additional concepts, specially that of codimension of a configuration. In section 5 we use the remarkable (albeit simple) connection between identically vanishing first moment (and thus the notion of equilibrium!) and the codimension of a configuration. This connection and corollary 4.2 produce the classical co-sphericity condition of Cayley [8]. The weight functions of systems with $\mu_{1}=\overrightarrow{O}$ (and fixed $X$) form a vector space $\mathbb{W}_{0}(X)$ of dimension equal to the codimension c:=nd1c:=n-d-1 of $X$, where $n$ is (throughout this paper) the number of points in $X$, and $d$ is dimension of the affine closure of $X$, usually referred to as the dimension of $X$. Theorem 4.1 implies a relation between the dimension of the configuration and vanishing Cayley-Menger determinants of subconfigurations of $X$ and, in the proof of theorem 5.3, we indicate how to obtain the independent (Cayley-Menger determinantal) constraints that ensure a configuration has a given dimension. Theorem 5.3 provides the expression $c+1\choose 2$} for the number of constraints that ensure a configuration has dimension d=nc1d=n-c-1. This expression generalizes the corresponding formulæ  for dimensions 1, 2 and 3 in Wintner [27], $\S 357$.
Section 6 concerns applications of the preceding sections to central configurations. We consider the not so explored problem of central configurations with zero total mass, for which some geometric properties and algebraic relations are directly determined. For central configurations with nonzero total mass, the fact that $\text{dim }\mathbb{W}_{0}(X)=c$ is used to extend to arbitrary codimensions some determinantal identities obtained by Dziobek for central configurations with codimension one [13, 16].
The article concludes with appendix A, where we apply the HLK theorem of section 2 to deduce simple linear constraints for squared mutual distances between points in space and the points of a fixed configuration.